Grade 11Mechanics

A particle is projected with a velocity 10 m/s at an angle 30 with horizontal . Find its distance from the point of projection 1/2 seconds later

Aayush
6 Years agoGrade 11

1 Answer

Vikas TU
6 Years ago
Dear student 
T = 2uSinQ /g 
T = 2*10 *1/20 = 1 sec
Displacement in 1/2 second 
X = uCosQ 
=> 10* sqrt 3 / 2 *1/2 = 4.33 meter 
Distance in y in t = 1/2 second 
Y = usin30 = 5 meter 
Distance = sqrt (4.33^2 + 25) = 6.614 m
Regards